Suppose you held a diversified portfolio consisting of a $7,500 investment in each of 20 different common stocks. The portfolio's beta is 1.47. Now suppose you decided to sell one of the stocks in your portfolio with a beta of 1.0 for $7,500 and use the proceeds to buy another stock with a beta of 2.27. What would your portfolio's new beta be? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Increase in beta as a result of net sales=2.27-1=1.27

Hence increase/stock=1.27/20=0.0635

Hence new beta=1.47+0.0635

=1.53(Approx).